OpenStudy (anonymous):

How do I integrate 1/2 ∫ cos (A+B) + cos (A-B) dx

OpenStudy (anonymous):

It depends on what A and B are.

OpenStudy (anonymous):

\[\int\limits_{}^{}cosAcosB\]

OpenStudy (anonymous):

Let just assume constant numbers and x only....

OpenStudy (anonymous):

Lets assume A is constant and B is combination of constant and x .

OpenStudy (anonymous):

\[1/2 ∫ cos (A+Bx) + cos (A-Bx) dx = 0.5(\frac{1}{B}sin(A+Bx)-\frac{1}{B}sin(A-Bx)+c)\]

OpenStudy (anonymous):

okay....i noe edy...thanks...

OpenStudy (anonymous):

\[\int cos(A)cos(Bx)dx = \frac{1}{B}cos(A)sin(Bx)\]